OpenStudy (anonymous):

Please help me... 2ln(x)+ln(x)^2=3

OpenStudy (anonymous):

I've gotten to this step lnx^2+lnx^2=3

OpenStudy (valpey):

So you can think of this as a quadratic equation where you are solving for ln(x). Let z = ln(x) \[z^2 + 2z -3 = 0\]

OpenStudy (valpey):

\[(z+3)*(z-1) = 0\] z = -3 or 1 ln(x) = -3 or 1 \[x = e^{-3} ,e^{1}\]

OpenStudy (valpey):

ln(x^2) isn't the same as (ln(x))^2

OpenStudy (valpey):

The equation editor is very useful in this kind of case. Do you know which one you mean? \[\ln{x^{2}}\] or \[(\ln{x})^{2}\]

OpenStudy (precal):

@amy_renae1994 Please tell us how you meant to post the problem. Use draw or equation editor

OpenStudy (anonymous):

\[2\ln x+\ln ^{2}=3\]

OpenStudy (precal):

what happened to the second x? can't have ln^2

OpenStudy (anonymous):

I'm sorry i meant to put it in there..

OpenStudy (anonymous):

Omg, I love you right now. I tried using wolframalpha earlier. Thanks so very much

OpenStudy (precal):

@Valpey I stand corrected @amy_renae1994 I erased my work so I do not confuse you

OpenStudy (valpey):

Bah, I totally forgot that the notation was: \[\ln^{2}{x}\]